An industrial setup can earn additional AED 7500 per year by installing a new production machine. If the applicable interest rate is 10% per year, how much should the company have over a 5 years to break even ?


A student spend $400, $700 and $ 600 in years 2, 4, and 7 on purchasing necessary stationery. The total worth of these spends in year 10 at an interest rate of 5% will be equivalent to:



How much money would be available in year 10 if $15000 is deposited each year in years 3 through 10 at an interest rate of 10% per year?
1
Expert's answer
2019-10-07T09:56:41-0400

"1) NPV = 7,500/1.1 + 7,500/1.1^2 + 7,500/1.1^3 + 7,500/1.1^4 + 7,500/1.1^5 = 28,430.9.
\\
2) NPV = 400/1.05^9 + 700/1.05^7 + 600/1.05^4 = 1,248.94.
\\
3) FV = 15,000×(1.1^7 - 1)/0.1 = 142,307.57."